# LLaMAX: Scaling Linguistic Horizons of LLM by Enhancing Translation Capabilities Beyond 100 Languages

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable translation capabilities in high-resource language tasks, yet their performance in low-resource languages is hindered by insufficient multilingual data during pre-training. To address this, we dedicate 35,000 A100-SXM4-80GB GPU hours in conducting extensive multilingual continual pre-training on the LLaMA series models, enabling translation support across more than 100 languages. Through a comprehensive analysis of training strategies, such as vocabulary expansion and data augmentation, we develop LLaMAX. Remarkably, without sacrificing its generalization ability, LLaMAX achieves significantly higher translation performance compared to existing open-source LLMs (by more than 10 spBLEU points) and performs on-par with specialized translation model (M2M-100-12B) on the Flores-101 benchmark. Extensive experiments indicate that LLaMAX can serve as a robust multilingual foundation model. The code 111https://github.com/CONE-MT/LLaMAX/. and models 222https://huggingface.co/LLaMAX/. are publicly available.
@@ 1 Introduction
Large Language Models (LLMs; Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; OpenAI, 2023; Touvron et al., 2023a, b) exhibit excellence in translation tasks involving high-resource languages (Vilar et al., 2023; Zhu et al., 2024b), yet their effectiveness in low-resource translation is suboptimal (Hendy et al., 2023; Bang et al., 2023; Zhu et al., 2024b). Figure 1 illustrates the number of translation directions with performance exceeding 10 spBLEU (Goyal et al., 2022) score on Flores-101 (Goyal et al., 2022). It is evident the majority of models are clustered around the origin point for Arabic-centric translations, demonstrating a significant disparity when compared to their English-centric performance.
This discrepancy is primarily due to the lack of pre-training data for these languages (Wei et al., 2023; Yuan et al., 2023b; Alves et al., 2024). Many researchers are actively working to address this issue. Guo et al. (2024) enhance the LLMs’ ability by translating low-resource languages after learning textbooks. Zhu et al. (2024b) find cross-lingual examples that can provide better task guidance for low-resource translation. In addition to the efforts focus on the fine-tuning stage, some studies have attempted to train a multilingual LLM from scratch (Wei et al., 2023), or to train a language-specific LLM (Faysse et al., 2024; Alves et al., 2024; Cui et al., 2024). However, the languages covered by these works are not extensive (Wei et al., 2023; Alves et al., 2024; Luo et al., 2023), and the translation performance is still unsatisfactory (Wei et al., 2023; Alves et al., 2024; Luo et al., 2023).
@@
A primary challenge in expanding language support lies in determining the appropriate vocabulary (Cui et al., 2024; Fujii et al., 2024). To face this, we conduct a quantitative analysis, assessing the impact of adding language-specific tokens from various angles: tokenization granularity, embedding quality, and the model’s inner distribution. Introducing a small number of new tokens significantly degrades existing LLM performance, while a larger new token set increases training complexity and data requirements. Surprisingly, adhering to the original vocabulary of LLMs emerges as the most cost-effective strategy for extending LLMs to 102 languages.
Another great challenge in extending language support is the scarcity of data for low-resource languages (Chang et al., 2023; Guo et al., 2024). To alleviate the scarcity of training data, we delve into dictionary-based data augmentation (Pan et al., 2021) and conduct a comprehensive analysis of various augmentation strategies. This analysis takes into consideration different dictionaries and data sources (monolingual or parallel data). We find that the optimal approach for data augmentation involves using parallel data, with the choice of dictionary correlated to the number of target language entities it covers.
@@
Comprehensive analysis of the key techniques in multilingual continual pre-training to LLMs, including vocabulary extension and data augmentation.
@@
2 Training Data Construction
2.1 Components of Training Data
During the continual pertaining stage, the collected training data covering 102 languages (refer to , which are all languages supported by Flores-101), mainly consists of two parts: monolingual () and parallel () data. For languages with limited data availability, we generated a pseudo-parallel dataset () with multilingual dictionaries: MUSE (Lample et al., 2018) and PanLex (Wang et al., 2022). More details regarding the supported languages, dataset description, and data statistics can be found in Appendix B.
Monolingual Data ().
Our monolingual training data includes 94 languages supported by Flores-101 from MC4 (Xue et al., 2021) and MADLAD (Kudugunta et al., 2024), totaling 40,000,000 sentences. To ensure efficient handling and processing of the data, we take a strategy where each piece of monolingual data is split into multiple entries, with a block size of 512.
Parallel Data ().
Our parallel data from Lego-MT Yuan et al. (2023a) encompasses 102 languages, forming a total language pairs and translation directions. For each translation direction, denoted as source language () to target language (), we concatenate each translation set, merely using a space as a delimiter, to form a single entry for training data. For each language pair, the probability of occurrence for each translation direction, for example, and is set as 50%. During the training stage, the gradient is computed for the entire data entry, rather than solely for the target sentence. For language pairs that have fewer than 25,000 (bound by machine resources) sentence pairs, we replicate the original data thrice (Muennighoff et al., 2023).
Data Generated Through Augmentation ().
The way to obtain code-switch data consists of two steps: 1) build multilingual lexicons; 2) construct pseudo-parallel data. We show the data augmentation process in Figure 2.
Step 1: Building multilingual lexicons.
[…]
@@ 2.1 Components of Training Data
Step 2: Constructing pseudo-parallel data.
We further conduct an experimental analysis in Section 3, and find the augmentation based on parallel data outperforms that on monolingual data. Therefore, the data augmentation is merely based on parallel data during continual pretraining.
2.2 Training Algorithm.
Given an LLM on a collected training data , where is the pre-trained parameters, our objective is to obtain an LLM through continual pre-training, denoted as . Here, indicates the updated parameters. The target of is to preserve the general capabilities of the model in high-resource languages while simultaneously enhancing the translation performance across all translation directions among 102 languages. The process of constructing training data is outlined in Algorithm 1. We gather monolingual data for each of the languages and parallel data for every translation direction. In particular, there is no augmentation for translations involving high-resource languages. Instead, we solely augment the translation data that is insufficient by utilizing a trained translation model, Lego-MT model. Then we train the , the loss function is calculated as:
where is the total decoding time step.
@@
3 Key Technique Design
In this section, we analyze primarily two key challenges related to the extension of language support: determining an appropriate vocabulary (in Section 3.1) and improving the effectiveness of data augmentation (in Section 3.2). For a more detailed analysis, refer to the discussions on the selection of multi-hop translation in the lexicon (see Appendix F) and the format of parallel data during continual pre-training (see Appendix G).
3.1 An Appropriate Vocabulary: The Original One.
Existing Pipeline.
Exploring adapting pre-trained LLMs to new languages without starting from scratch seems to have a concise pipeline, resulting in ChineseLLaMA2 (Cui et al., 2024), Swallow (Fujii et al., 2024), and so on. This pipeline comprises three crucial steps: 1) vocabulary expansion: extending the vocabulary of LLMs by adding new tokens specific to that language and initializing these new tokens as the average of embeddings from the existing tokens (Dobler and de Melo, 2023). 2) continual pre-training: continual pre-training the LLM on a large corpus of text data from the target language. 3) instruction tuning: aligning the model with specific tasks or instructions, enhancing its performance. Instead of simply following the pipeline, we question the necessity of vocabulary expansion.
@@ 3 Key Technique Design
We conduct a series of analytical experiments on the LLaMA2 vocabulary. Our initial focus is on examining the correlation between fertility and the quality of token representation. Here, fertility refers to the ratio of the length of the token sequence produced by the LLaMA2 tokenizer to the length of the input sentence when split by spaces (Chinese and Japanese is split by character). Furthermore, we carry out experiments using 10,000 enro and enbn bilingual sentence pairs from Lego-MT dataset. In each experiment, we introduce a varying number of language-specific new tokens and evaluate each model on the Flores-101.
@@
We assess the quality of representation by enX translation task. This task identifies the translated result that best aligns with the corresponding English sentence within an extensive target dataset, and evaluates with Recall at top 1, denoted as R@1 (Kabir and Carpuat, 2021). A higher R@1 value signifies a more robust quality of the representation. Concurrently, we present the cosine similarity of representations generated by LLaMA2 for identical sentences in English and other languages. On experiments across 102 languages, more details in Appendix D, there exists a strong correlation between fertility and the quality of representation, evidenced by a Spearman correlation coefficient of approximately -0.88 for each assessed quality metric.
@@
As demonstrated in Table 1, even the addition of a small number (100) of new language-specific tokens can have a significant impact on the multilingual performance of LLMs. In addition, we conduct a further analysis on the original tokens (32k) embedding distribution and the token number before and after adding new tokens by KS-Lottery (Yuan et al., 2024). For more details on KS-Lottery, refer to Appendix E. As the experimental result of “shift distance” and “# shift token” in Tabel 1, fine-tuning the entire model with limited new tokens follows a similar pattern to that with the original vocabulary. However, an excessive number of new tokens can shift the model’s training focus. This holds true regardless of whether the language (ro) is well-supported by the model or not (bn). The influence of these additional tokens is substantial, indicating that the process of enhancing the multilingual capabilities of LLMs is not as straightforward as simply expanding the vocabulary and training with more multilingual data.
| enX | Xen | taX | Xta | thX | Xth | zhX | Xzh | ||
| LLaMA2 | ✗ | 18.31 | 23.61 | 0.99 | 0.49 | 4.83 | 1.15 | 10.02 | 7.35 |
| ✗ | 19.06 | 25.98 | 3.20 | 0.91 | 7.66 | 3.13 | 11.32 | 7.83 | |
| + | ✗ | 19.46 | 26.40 | 4.17 | 1.76 | 7.28 | 3.02 | 11.65 | 8.82 |
| + | ✗ | 19.22 | 25.91 | 3.51 | 1.34 | 7.64 | 2.83 | 11.56 | 7.99 |
| ++ | ✗ | 19.36 | 26.47 | 4.35 | 1.82 | 7.78 | 3.49 | 11.44 | 9.14 |
| + | ✓ | 19.47 | 26.65 | 4.54 | 1.83 | 7.66 | 3.13 | 11.89 | 9.17 |
| + | ✓ | 18.59 | 25.98 | 3.61 | 1.36 | 6.72 | 2.35 | 10.81 | 6.45 |
| ++ | ✓ | 19.70 | 26.71 | 4.68 | 1.82 | 8.21 | 3.65 | 12.05 | 9.28 |
| ++ | ✓ | 19.17 | 26.58 | 4.57 | 1.95 | 7.12 | 3.12 | 11.52 | 7.73 |
| ++ | ✓ | 18.80 | 26.56 | 4.78 | 1.79 | 7.31 | 3.18 | 11.35 | 7.28 |
@@
The LLaMA tokenizer, which utilizes the Byte-level Byte Pair Encoding (BBPE; Wang et al., 2019) algorithm, is the foundation for multilingual language processing tasks. Its universal compatibility across all languages, in conjunction with the absence of the requirement for an “unknown” token, optimizes vocabulary sharing (Yuan et al., 2023b) and improves its robustness. It allows the model to understand/generate responses in various languages using the same vocabulary. Meanwhile, studies have shown that LLMs trained on unbalanced English-centric datasets, often use English as an internal pivot language. This helps LLMs to map the inputs closer to English in internal space before generating the output (Zhu et al., 2024a; Yoon et al., 2024). Maintaining the original vocabulary helps to preserve this behavior, which also benefits for improving the multilingual capability.
@@ 3.2 Data Augmentation
Finding: The choice of dictionary is related to the number of entities for the language in the dictionary.
4 Benchmarking Results
In this section, we present multilingual benchmarking results to comprehensively demonstrate the potential of LLaMAX2. We evaluate translation quality with spBLEU (Goyal et al., 2022) and COMET-22 (Rei et al., 2020) for both LLMs and translation models. See Appendix C for training details on LLaMAX2 and description of baseline models.
@@ 4 Benchmarking Results
Next, we compare our language-extended LLaMAX2-Alpaca model with other open-source decoder-only LLMs built for multilingual purposes (Table 4, Table 5). Compared to other from-scratch trained LLMs, such as PolyLM, Yayi2, LLaMAX2 consistently shows better performance across various multilingual translation benchmarks, indicating that the LLaMA2 base model provides a strong foundation for language extension. Furthermore, when compared to other LLaMA-based continual pre-trained models, such as TowerInstruct, LLaMAX2 also achieves superior performance, demonstrating the effectiveness of our optimized continual pre-training pipeline.

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable translation capabilities in high-resource language tasks, yet their performance in low-resource languages is hindered by insufficient multilingual data during pre-training. To address this, we conduct extensive multilingual continual pre-training on the LLaMA series models, enabling translation support across more than 100 languages. Through a comprehensive analysis of training strategies, such as vocabulary expansion and data augmentation, we develop LLaMAX. Remarkably, without sacrificing its generalization ability, LLaMAX achieves significantly higher translation performance compared to existing open-source LLMs (by more than 10 spBLEU points) and performs on-par with specialized translation model (M2M-100-12B) on the Flores-101 benchmark. Extensive experiments indicate that LLaMAX can serve as a robust multilingual foundation model. The code 111https://github.com/CONE-MT/LLaMAX/. and the models 222https://huggingface.co/LLaMAX/. are publicly available.
@@ 1 Introduction
Large Language Models (LLMs; Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; OpenAI, 2023; Touvron et al., 2023a, b) exhibit excellence performance in translation tasks involving high-resource languages (Vilar et al., 2023; Zhu et al., 2024b), yet their effectiveness in low-resource translation is suboptimal (Hendy et al., 2023; Bang et al., 2023; Zhu et al., 2024b). Figure 1 illustrates the number of translation directions with performance exceeding 10 spBLEU (Goyal et al., 2022) score on Flores-101 (Goyal et al., 2022). It is evident the majority of models are clustered around the origin point for Arabic-centric translations, demonstrating a significant disparity when compared to their English-centric performance.
This discrepancy is primarily due to the lack of pre-training data for these languages (Wei et al., 2023; Yuan et al., 2024b; Alves et al., 2024). Many researchers are actively working to address this issue. Guo et al. (2024) enhance the LLMs’ ability by translating low-resource languages after learning textbooks. Zhu et al. (2024b) find cross-lingual examples that can provide better task guidance for low-resource translation. In addition to the efforts focus on the fine-tuning stage, some studies have attempted to train a multilingual LLM from scratch (Wei et al., 2023), or to train a language-specific LLM (Faysse et al., 2024; Alves et al., 2024; Cui et al., 2024). However, the languages covered by these works are not extensive (Wei et al., 2023; Alves et al., 2024; Luo et al., 2023), and the translation performance is still unsatisfactory (Wei et al., 2023; Alves et al., 2024; Luo et al., 2023).
@@
A primary challenge in expanding language support lies in determining the appropriate vocabulary (Cui et al., 2024; Fujii et al., 2024). After assessing the impact of adding language-specific tokens from various angles: tokenization granularity, embedding quality, and the model’s inner distribution, we find that introducing a small number of new tokens significantly degrades existing LLM performance, while a larger new token set increases training complexity and data requirements. Surprisingly, adhering to the original vocabulary of LLMs emerges as the most cost-effective strategy for extending LLMs to 102 languages.
Another great challenge in extending language support is the scarcity of data for low-resource languages (Chang et al., 2023; Guo et al., 2024). To alleviate the scarcity of training data, we delve into dictionary-based data augmentation (Pan et al., 2021; Lu et al., 2023) and conduct a comprehensive analysis of various augmentation strategies. This analysis takes into consideration different dictionaries and data sources (monolingual or parallel data). We find that the optimal approach for data augmentation involves using parallel data, with the choice of dictionary correlated to the number of target language entities it covers.
@@
Comprehensive analysis of the key techniques in multilingual continual pre-training, including vocabulary extension and data augmentation.
@@
2 Key Technique Design
@@
Existing Pipeline.
Exploring adapting pre-trained LLMs to new languages without starting from scratch seems to have a concise pipeline, resulting in ChineseLLaMA2 (Cui et al., 2024), Swallow (Fujii et al., 2024), and so on. This pipeline comprises three crucial steps: 1) vocabulary expansion: extending the vocabulary of LLMs by adding new tokens specific to that language and initializing these new tokens as the average of embeddings from the existing tokens (Dobler and de Melo, 2023). 2) continual pre-training: continual pre-training LLM on a large corpus of text data from the target language. 3) instruction tuning: aligning the model with specific tasks or instructions, enhancing its performance. Instead of simply following the pipeline, we analyze primarily two key challenges related to the extension of language support: determining an appropriate vocabulary (in Section 2.1) and improving the effectiveness of data augmentation (in Section 2.2). For a more detailed analysis, refer to the discussions on the selection of multi-hop translation in the lexicon (see Appendix E) and the format of parallel data during continual pre-training (see Appendix F).
@@
2.1 Existing Vocabulary is Adequate.
@@
We conduct a series of analytical experiments on the LLaMA2 vocabulary. Our initial focus is on examining the correlation between fertility and the quality of token representation. Here, fertility refers to the ratio of the length of the token sequence produced by the LLaMA2 tokenizer to the length of the input sentence when split by spaces (Chinese and Japanese is split by character). Furthermore, we carry out experiments using 10,000 enro and enbn bilingual sentence pairs from Lego-MT dataset. For new tokens, the BBPE algorithm is executed on language-specific data from MADLAD-400 to produce a vocabulary of 100,000 tokens. Within this vocabulary, language-specific tokens are arranged based on their frequency in the corpus. Subsequently, we identify the top-k tokens (where k is determined by the corresponding “#New Token” in Table 1) that are absent in the original LLaMA vocabulary and incorporate them as new tokens into the LLaMA vocabulary. In each experiment, we introduce a varying number of language-specific new tokens and evaluate each model on the Flores-101.
@@
We assess the quality of representation by enX translation task. This task identifies the translated result that best aligns with the corresponding English sentence within an extensive target dataset, and evaluates with Recall at top 1, denoted as R@1 (Kabir and Carpuat, 2021). A higher R@1 value signifies a more robust quality of the representation. Concurrently, we present the cosine similarity of representations generated by LLaMA2 for identical sentences in English and other languages. On experiments across 102 languages, more details in Appendix C, there exists a strong correlation between fertility and the quality of representation, evidenced by a Spearman correlation coefficient of approximately -0.88 for each assessed quality metric.
@@
As demonstrated in Table 1, even the addition of a small number (100) of new language-specific tokens can have a significant impact on the multilingual performance of LLMs. In addition, we conduct a further analysis on the original tokens (32k) embedding distribution and the token number before and after adding new tokens by KS-Lottery (Yuan et al., 2024a). For more details on KS-Lottery, refer to Appendix D. As the experimental result of “shift distance” and “# shift token” in Tabel 1, fine-tuning the entire model with limited new tokens follows a similar pattern to that with the original vocabulary. However, an excessive number of new tokens can shift the model’s training focus. This holds true regardless of whether the language (ro) is well-supported by the model or not (bn). The influence of these additional tokens is substantial, indicating that the process of enhancing the multilingual capabilities of LLMs is not as straightforward as simply expanding the vocabulary and training with more multilingual data.
@@
The LLaMA tokenizer, which utilizes the Byte-level Byte Pair Encoding (BBPE; Wang et al., 2019) algorithm, is the foundation for multilingual language processing tasks. Its universal compatibility across all languages, in conjunction with the absence of the requirement for an “unknown” token, optimizes vocabulary sharing (Yuan et al., 2024b) and improves its robustness. It allows the model to understand/generate responses in various languages using the same vocabulary. Meanwhile, studies have shown that LLMs trained on unbalanced English-centric datasets, often use English as an internal pivot language. This helps LLMs to map the inputs closer to English in internal space before generating the output (Zhu et al., 2024a; Huang et al., 2024b; Yoon et al., 2024). Maintaining the original vocabulary helps to preserve this behavior, which also benefits for improving the multilingual capability.
@@ 2.2 Data Augmentation
Finding: The choice of dictionary is related to the number of entities in the dictionary.
3 Training Data Construction
3.1 Components of Training Data
During the continual pertaining stage, the collected training data covering 102 languages (refer to , which are all languages supported by Flores-101), mainly consists of two parts: monolingual () and parallel () data. For languages with limited data availability, we generated a pseudo-parallel dataset () with multilingual dictionaries: MUSE (Lample et al., 2018) and PanLex (Wang et al., 2022). The whole continual pre-training utilizes over 64 billion tokens. More details on supported languages, dataset description, and data statistics can be found in the Appendix A.
Monolingual Data ().
Our monolingual training data includes 94 languages supported by Flores-101 from MC4 (Xue et al., 2021) and MADLAD (Kudugunta et al., 2024), totaling 40,000,000 sentences. To ensure efficient handling and processing of the data, we use a strategy in which each piece of monolingual data is split into multiple entries, with a block size of 512.
Parallel Data ().
Our parallel data from Lego-MT Yuan et al. (2023) encompasses 102 languages, forming a total language pairs and translation directions. For each translation direction, denoted as source language () to target language (), we concatenate each translation pair, merely using a space as a delimiter, to form a single entry for training data. For each language pair, the probability of occurrence for each translation direction, for example, and is set as 50%. During the training stage, the gradient is computed for the entire data entry, rather than only for the target sentence. For language pairs that have fewer than 25,000 (bound by machine resources) sentence pairs, we replicate the original data thrice (Muennighoff et al., 2023).
Data Generated Through Augmentation ().
The way which is followed by Pan et al. (2021), to obtain code-switch data consists of two steps: 1) build multilingual lexicons; 2) construct pseudo-parallel data. We show the data augmentation process in Figure 2.
@@ 3.1 Components of Training Data
Step 1: Building multilingual lexicons.
[…]
Step 2: Constructing pseudo-parallel data.
3.2 Training Algorithm.
Given an LLM on a collected training data , where is the pre-trained parameters, our objective is to obtain an LLM through continual pre-training, denoted as . Here, indicates the updated parameters. The target of is to preserve the general capabilities of the model in high-resource languages while simultaneously enhancing the translation performance across all translation directions among 102 languages. The process of constructing training data is outlined in Algorithm 1. We gather monolingual data for each of the languages and parallel data for every translation direction. In particular, there is no augmentation for translations involving high-resource languages. Instead, we solely augment the translation data that is insufficient by utilizing a trained translation model, Lego-MT model. Then we train the , the loss function is:
where is the total decoding time step.
4 Benchmarking Results
In this section, we present multilingual benchmarking results to comprehensively demonstrate the potential of LLaMAX2. We evaluate translation quality with spBLEU (Goyal et al., 2022) and COMET-22 (Rei et al., 2020) for both LLMs and translation models. See Appendix B for training details on LLaMAX2 and description of baseline models.
@@ 4 Benchmarking Results
Next, we compare LLaMAX2-Alpaca model with other open-source decoder-only LLMs built for multilingual purposes (Table 3, Table 4). Compared to other from-scratch trained LLMs, such as PolyLM, Yayi2, LLaMAX2 consistently shows better performance across various multilingual translation benchmarks, indicating that the LLaMA2 base model provides a strong foundation for language extension. Furthermore, when compared to other LLaMA-based continual pre-trained models, such as TowerInstruct, LLaMAX2 also achieves superior performance, demonstrating the effectiveness of our optimized continual pre-training pipeline.